Let f(x)=3x+2. Find f(f(1))
Answer by ikleyn(52780)

.

First calculate f(1) by substituting the value of 1 instead of x into the formula for f(x)


    f(1) = 3*1 + 2 = 3 + 2 = 5.


To find f(f(1)),  replace the interior f(1) by 5, since you just calculated it. 
You will get


    f(f(1)) = f(5) = 3*5 + 2 = 15 + 2 = 17.


ANSWER.  f(f(1)) = 17, and you know now how to find it.

Solved and explained.